OpenStudy (anonymous):

Given the points A(-3,-8) and B(3,-2), find the points P,(1,b) and P,,(c, d) which divides AB internally and externally in the same ratio.

OpenStudy (amistre64):

focus on one direction at a time x: -3, 1, c, 3 or -3, c, 1, 3 define c that fits the requirements,

OpenStudy (amistre64):

i do not know wht it means to be internally and externally ratioed so that part is up to you to figure out.

OpenStudy (mathmate):

Start with a sketch of the situation  As @amistre64 said, concentrate on one direction, say x. Given A(-3,-8), B(3,-2) and P(1,b) Find the ratio that P divides AB internally AP/PB = (1-(-3))/(3-1)=4/2=2 In the y-direction, the same ratio applies, so AP/PB=2=(b-(-8))/(-2-b) solve for B. Hint: it is a negaive integer. For external division, P'B is negative (see sketch). Post if you have questions.

OpenStudy (anonymous):

so b=-4?